OpenStudy (anonymous):

please help? Consider the function e^-8x-x^2-10y-y^2 Find fx,fy,fxx,fyy,fxy? i found fx=-2 e^(-8 x-x^2-y (10+y)) (4+x) and fy=-2 e^(-8 x-x^2-y (10+y)) (5+y) please help find the others? thanks :)

OpenStudy (abb0t):

By \(f_{xx}\) I assume you're referring to partial derivatives, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i cant figure it out! any help?

OpenStudy (anonymous):

i tried fxy as -2 e^(-8 x-x^2-y (10+y)) (4+x) (5+y) but didnt work! and fxx as 4 e^(-16 x-2 x^2-2 y (10+y)) (4+x)^2

OpenStudy (anonymous):

hey, partial derivative means what? it's not that. when you take derivative respect to x, y = constant. your post is not clear, is it \[e^-8x -x^2 -10y -y^2\]

OpenStudy (anonymous):

let F = your function, Fx = (e^-8x)' -(x^2)' only,

OpenStudy (anonymous):

sorry repost yours. I confused

OpenStudy (anonymous):

oh,, i see. ok nothing is big

OpenStudy (anonymous):

now Fx= ? hihihi,, so funny. I figure out that is a real big. sorry

OpenStudy (anonymous):

mine is Fx = (-8-2x) * the whole original

OpenStudy (anonymous):

is it yours?

OpenStudy (anonymous):

ok, check it please, i don't think mine is wrong. (e^u)' = u' e^u is it right?

OpenStudy (anonymous):

ok, now again, Fxx= take quotient rule for the whole thing = ? do it , we compete.

OpenStudy (anonymous):

I don't say compare, I say compete . play?

OpenStudy (anonymous):

you are loser!!!!!

OpenStudy (anonymous):

hey friend, this time, you have (-8-2x)' *whole thing + (wholething)' * (-8-2x) are you ok then?

OpenStudy (anonymous):

not digest it yet? you have to take derivative of a product

OpenStudy (anonymous):

to get Fxx, you have to take (Fx))' respect to x again, right? so you start at Fx = (-8-2x)*wholething, right?

OpenStudy (anonymous):

is it not take derivative of a product? is it not apply product rule?

OpenStudy (anonymous):

ok, you cannot see I cheat on you, I deleted that answer when I figure out that is not correct. hihihi,,,, learn something?

OpenStudy (anonymous):

that is just the first part, wait I put it in neat.

OpenStudy (anonymous):

\[F(x) = (-8-2x) *wholething\] Fxx = (-8-2x)' *wholething + (wholething)' ((-8-2x) = -2*wholething + (-8-2x) Fx (remember wholething ' =Fx?) = -2 *wholething + (-8-2x)(-8-2x)* wholething factor out = wholething* [(-2 + (-8-2x)^2) expand Fxx = wholething * (4x^2 +16x+66)

OpenStudy (anonymous):

no, stop there, when we are wrong, you are better check, recheck to figure out why we are wrong than go over it. I don't think it's good. ask other for help. I want to know why i am wrong, too. actually, I am not sure whether I can do anything else or not. I am better sit aside and look

OpenStudy (anonymous):

shame on us, right? 2 students cannot solve the problem.

OpenStudy (anonymous):

don't tell me you type a wrong answer to computer, so it says wrong. you cheat on me back???? ok, fair enough

OpenStudy (anonymous):

ok, but mine is wrong

OpenStudy (anonymous):

thanks for the link, i go there now. good luck, friend. I am a loser!!!!

OpenStudy (anonymous):

do it by yourself. I am in bad mood, cannot step any step. but I will be back later to know the right answer, don't close it, ok?

OpenStudy (anonymous):

it will only let me delete mine?

OpenStudy (anonymous):

ok will do :)

OpenStudy (anonymous):

I got it, thank you very much. close it now.

OpenStudy (anonymous):

ok :)